OpenStudy (anonymous):

solve by factoring show all work thank you :D x^2+2x-3=0

OpenStudy (anonymous):

you take the formula, bust it up into this: (x+3)(x-1) Now, you use the zero product property, which allows you to divide one out, placing the other equal to zero. So, you should have two equations: (x+3)=0 (x-1)=0 Now, simply solve for x x=-3, x=1

OpenStudy (anonymous):

(x+3)(x-1)